OpenStudy (anonymous):

siplify 12w ---- 4w^4

OpenStudy (anonymous):

divide both by 4

OpenStudy (anonymous):

so 3w over w^4

OpenStudy (anonymous):

right, now divide by w

OpenStudy (anonymous):

how the heck do you divid by w?

OpenStudy (anonymous):

Assuming\[12\times w \div (4\times w ^{4}) =3\div w ^{3}\] or in another way \[3\times w ^{-3}\] Cheers

OpenStudy (anonymous):

w is just representing a number, so, because w is on both levels, you can divide by it. so you have this right now,\[3w/w ^{4}\] which is just \[3\times w/w \times w \times w \times w\] You can cancel (AKA divide) the w.

OpenStudy (anonymous):

So can you tell me what you'd end up with?

OpenStudy (anonymous):

3 over w^3

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

So do you understand how we got that answer?

OpenStudy (anonymous):

nope..

OpenStudy (anonymous):

lol okay, do you have a specific question?

OpenStudy (anonymous):

or is it just the concept in general?

OpenStudy (anonymous):

nope i just dont get any of it

OpenStudy (anonymous):

alright, well when you simplify a fraction, you can factor them. All factoring is is breaking it down into little numbers, so your equation after it's been factored would be this....\[3\times4\times w/4 (w \times w \times w \times w)\]

OpenStudy (anonymous):

Get that? So if you did all that math, you'd end up with your original equation.

OpenStudy (anonymous):

The next step is to see what numbers (or letters) the upper and lower levels have in common. So the only numbers/letters that are in common are 4 and w.

OpenStudy (anonymous):

Because they're in common (and they're both multiplied) you can cancel them. All canceling is is doing the OPPOSITE of whats being done. So, because they're being multiplied, We would DIVIDE to get ride of them.

OpenStudy (anonymous):

thats the second step i took to get rid of 4.

OpenStudy (anonymous):

AND also the third step I took to get rid of w.

OpenStudy (anonymous):

So if we put a number in for w, just so you can see it with real numbers, We'd have this. \[3\times 4\times2/4(2\times2\times2\times2)\] remember that 2 is just a random number i picked to represent w. So now can you see better how we canceled w?